OpenStudy (anonymous):

Four of the angles of a pentagon measure 85°, 110°, 135°, and 95°. Find the measure of the missing angle. 115° 95° 135° 85°

OpenStudy (anonymous):

Max angles of a polygon are given by: 180(n-2) where n is number of sides: So: for a pentagon = 5 sides. 180(5-2) = 180*3 = 540 degrees Then, subtract the four angles and you get: 540 - 110 - 95 - 85 -135 = 115 degrees for missing angle.

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

No problem kiddo.